INFRARED SPECTRUM OF THE (CO_{2})_{2}-N_{2}O TRIMER MEASURED IN N_{2}O ν_{1} AND ν_{3} REGIONS.

\underline{J. NOROOZ OLIAEE}, N. MOAZZEN-AHMADI, Department of Physics and Astronomy, University of Calgary, 2500 University Dr., N.W., Calgary, Alberta T2N 1N4, Canada; A.R.W. McKELLAR, National Research Council of Canada, Ottawa, Ontario, K1A 0R6, Canada.

Infrared spectra of the (CO_{2})_{2}-N_{2}O trimer are observed by exciting the ν_{1} and ν_{3} fundamental stretching vibrations of the N_{2}O moiety (around 1285 cm^{−1} and 2224 cm^{−1} respectively). Spectra are recorded using a pulsed supersonic jet apparatus with a tunable diode laser probe. Ground state parameters were previously determined from a microwave study.^{a} Analysis of the infrared spectra reveals information on the vibrational shifts upon complex formation as well as molecular parameters for the excited states. Our cluster calculation program yields a minimum energy structure very similar to that from the Orient program^{a} but in slightly better agreement with the experimental structure. Our cluster calculations indicate a close resemblance of the two lowest energy isomers to those of (CO_{2})_{2}-OCS.

aR. A. Peebles, S. A. Peebles, and R. L. Kuczkowski Mol. Phys. 96, 1355 (1999).